OpenStudy (anonymous):

Train A and B are traveling in the same direction on parael tracks. Train A is traveling at 80 miles per hour and Train B traveling at 100 miles per hour. Train A passes a station at 6:15 pm. If train B passes the same staton at 6:27pm what time will train b catch up t6o train a

OpenStudy (anonymous):

Train A=80x+12*80 Train B= 100x 100x=80x+12*80 x=48

OpenStudy (anonymous):

6:27+ 48 minutes

OpenStudy (anonymous):

7:15

OpenStudy (anonymous):

Since there is 12 minutes between the 2 trains passing the station, at the time train B passes the station, train A has traveled 16 miles (1/5 of 80). Train B is making up 20 miles of the difference every hour. So in 16/20 = 4/5 of an hour, train B will catch up to train A. 4/5 of 60 minutes is 48 minutes. So, 6:27 + 0:48 = 7:15pm